OpenStudy (anonymous):

Part A: Nadia rented a stage platform at $410 for 4 days. If she rents the same stage platform for 8 days, she has to pay a total rent of $770. Write an equation in the standard form to represent the total rent (y) that Nadia has to pay for renting the stage platform for x days.

OpenStudy (anonymous):

Thank you so much for helping me out. :D

OpenStudy (superdavesuper):

Standard form equation is y=Ax + B where A and B are numbers. So u have to calculate them. This is done by using the two cases: 410 = A*(4) + B 770 = A*(8) + B Subtract the two equations and u will find A; then calculate B to get ur answer :)

OpenStudy (anonymous):

a represents b represents

OpenStudy (superdavesuper):

Just two fixed numbers in ur standard equation but u have to find them in this case.

OpenStudy (anonymous):

ok can you help me with two other related questions?

OpenStudy (superdavesuper):

Sure i can try :)

OpenStudy (anonymous):

Write the equation obtained in Part A using function notation.

OpenStudy (anonymous):

Part a is what i first posted

OpenStudy (superdavesuper):

Hehee u need 2 calculate A and B to get the standard equation first... Then the function notation is easy!

OpenStudy (anonymous):

ya btw i really dont understand how to find a or b

OpenStudy (anonymous):

u said to subtract the equations but wat part

OpenStudy (superdavesuper):

The whole parts - both sides of the = sign. Try it n tell me wat u find.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

770 - 410= 360 8 - 4= 4

OpenStudy (superdavesuper):

OK good job @Polkadotgamer :) So after subtraction, u get: 360 = 4A right? What is the value of A then?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

360/4=90

OpenStudy (superdavesuper):

heheee almost there - what is B then?

OpenStudy (anonymous):

no clue i think 4

OpenStudy (superdavesuper):

Put A=90 back in the first equation, what is B?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

for both equations

OpenStudy (anonymous):

410 = 90*(4) + B B = 50 770 = 90*(8) + B B = 50

OpenStudy (superdavesuper):

VERY GOOD :) So the standard equation is what now?

OpenStudy (anonymous):

so how about the function notation

OpenStudy (superdavesuper):

Gimme the standard eqn first plz? :D

OpenStudy (anonymous):

410 = 90*(4) + B B = 50 770 = 90*(8) + B B = 50

OpenStudy (anonymous):

410 = 90*(4) + 50 770 = 90*(8) + 50

OpenStudy (superdavesuper):

Hmmm....nope....the standard equation that shows x and y....

OpenStudy (superdavesuper):

There should only be one....

OpenStudy (anonymous):

i thought we could do both but i guess 410 = 90*(4) + 50

OpenStudy (superdavesuper):

But where is x and y in ur eqn?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

rent (y) that Nadia has to pay for renting the stage platform for x days

OpenStudy (superdavesuper):

yup yup so keep x and y but replace A and B with what u have calculated....

OpenStudy (anonymous):

so 410 is y and x is 4

OpenStudy (superdavesuper):

Nope....hmmm....do u know what is variable and what is constant in an equation?

OpenStudy (anonymous):

a variable changes and a constant stays the same

OpenStudy (anonymous):

the variable is 4?

OpenStudy (superdavesuper):

yup yup so 4 cannot be a variable cuz' it does not change....x on the other hand is a variable. got it?

OpenStudy (anonymous):

yes so after that?

OpenStudy (superdavesuper):

So write out the equation with x, y as the variables and A, B as constants from wat u have calculated...

OpenStudy (anonymous):

a = x*(b) + y

OpenStudy (superdavesuper):

Go back to my first post and see what the equation looks like, then replace A, B with ur answers

OpenStudy (anonymous):

thats what i did

OpenStudy (superdavesuper):

The equation is y=Ax + B.....

OpenStudy (anonymous):

omgosh

OpenStudy (superdavesuper):

got it? :D :D :D

OpenStudy (anonymous):

so thats the answer to the first question and all u do for the second question is replace y with f of x

OpenStudy (superdavesuper):

Congrats @Polkadotgamer u just solve ur prob :D

OpenStudy (anonymous):

omygosh

OpenStudy (superdavesuper):

Wait wait wait show me the standard equation 1st plz?

OpenStudy (anonymous):

y=Ax + B